# From Words to Collisions: LLM-Guided Evaluation and Adversarial Generation of Safety-Critical Driving Scenarios

### v3

@@ II-A Safety-Critical Driving Scenario Evaluation for ADS
LLMs as evaluator: Recent studies have explored LLMs for driving scenario evaluation, focusing on realism, behavior, and consistency. In [13], they prompted LLMs with contextual data, e.g., road layout, weather, vehicle behavior, to assess scenario realism using the DeepScenario dataset [14]. [15] proposes a framework incorporating high-level evaluation criteria, including performance, safety, and comfort, alongside scenario data from the CARLA [16] to evaluate driving style and performance. However, in [15] they focus on assessing final driving style and driving performance level based on aggregated input, rather than directly assessing individual metrics like safety. OmniTester [17] extended text-only evaluation to multimodal analysis by generating captions from visualized scenarios using a vision-language model and comparing them with original descriptions from LLM.
@@ III-C2 Prompt Formulations
The safety-critical metrics prompt follows the structure of $\textit{Prompt}_{\mathrm{cart}}$ and $\textit{Prompt}_{\mathrm{ego}}$, with the LLM acting as a domain expert in collision evaluation. It incorporates CoT reasoning and ICL examples to guide the assessment of overall risk scores for each obstacle, based on thresholds for each safety-critical metric. The prompt also includes motion descriptions, as shown in Fig. 5(a), focusing on obstacles approaching from the front or sides, since these directions are more relevant for potential collision scenarios crucial to evaluating planner performance. This enhanced structure enables the LLM to systematically reason about ego-attackers using both quantitative metrics and spatial context.
For trajectory generation, we adopt a different prompting strategy as the LLM transitions from evaluator to generator. The system message instructs the model to synthesize an adversarial trajectory that increases the collision likelihood with the EV, given the scenario context. To enhance the generation performance, we apply Self-Consistency (SC) by prompting the model to produce multiple adversarial trajectory candidates for the same ego-attacker. Each candidate is evaluated using the safety-critical metrics prompt, and the trajectory with the lowest (i.e., most critical) risk score for the ego-attacker is selected.
@@ III-D Structured Output Format
In our LLM-based scenario evaluation and generation framework, we enforce structured output formats using predefined schemes, which ensure consistency and reduce hallucinated outputs [33]. These formats define fixed keys aligned with specific use cases, including collision evaluation, safety-critical scenario inference, agent-level risk assessment (e.g., risk scores), and trajectory generation parameters (e.g., position, velocity). The structure remains flexible and can be adapted to various prompt designs and task requirements.
@@ IV-B Collision Evaluation in Driving Scenarios
This use case evaluates if collisions occur in driving scenarios with EV. The inputs of our LLM model include a comprehensive map with geometric lanelet information, dynamics status of the EV, and of all traffic participants throughout the simulation’s duration. Our model outputs a boolean flag to indicate if a collision occurs for the EV in the given scenario. We randomly select 100 non-collision and 100 collision scenarios for evaluation from the 6000 total scenarios.
@@ IV-B1 Ablation Study
The performance of our LLM-based evaluator depends on prompt design and advanced techniques with Contextual Prompting (CP), In-Context Learning (ICL), and Chain-of-Thought (CoT) reasoning, as described in Section III-B2. To evaluate their impact, we conduct an ablation study comparing four configurations for both $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$: Base (no techniques), CP (CP only), CoT (CP + CoT), and All (CP + CoT + ICL).
@@ IV-B2 Comparison Across Different LLMs
We compare the reasoning performance of Gemini-1.5Pro, GPT-4o, and DeepSeek-V3 using the same $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ with All configuration applied. As shown in Table II, all models perform significantly better with $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$, reinforcing the findings in Section IV-B1. Gemini-1.5Pro achieves the highest F1 score (0.21) with $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$, likely due to its larger context window (up to 200k tokens), which helps interpret verbose Cartesian inputs. With ego-centric prompts, DeepSeek-V3 with an F1 score of 0.90 slightly outperforms the others, followed by GPT-4o (0.88) and Gemini-1.5Pro (0.87), suggesting a slight edge in reasoning ability for DeepSeek-V3.
@@ IV-D Generation of Safety-Critical Scenarios
Figure 8: Risk score of the ego-attacker vehicle’s trajectory, in the original safe scenario and in the generated safety-critical one. A risk score 5 indicates very low risk, while 0 denotes a collision.
@@ V Conclusion & Outlook
The framework was evaluated in three use cases supported by structured outputs: collision scenario evaluation, safety inference, and safety-critical scenario generation. Across all LLMs, $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ consistently outperformed $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$, for the Gemini-1.5Pro, achieving F1 scores of 0.87 vs. 0.21 in collision evaluation, and 0.83 vs. 0.62 in safety inference, highlighting the advantage of ego-centric representations for reasoning about risk. In generation, the framework successfully minimized the ego-attacker’s risk score to synthesize safety-critical outcomes. Future work includes extending the framework to 3D simulators such as CARLA, and enhancing the generation module to produce more diverse and challenging safety-critical scenarios.

### v4

@@ II-A Safety-Critical Driving Scenario Evaluation for ADS
LLMs as evaluator: Recent studies have explored LLMs for driving scenario evaluation, focusing on realism, behavior, and consistency. In [12], they prompted LLMs with contextual data, e.g., road layout, weather, vehicle behavior, to assess scenario realism using the DeepScenario dataset [13]. [14] proposes a framework incorporating high-level evaluation criteria, including performance, safety, and comfort, alongside scenario data from the CARLA [15] to evaluate driving style and performance. However, in [14] they focus on assessing final driving style and driving performance level based on aggregated input, rather than directly assessing individual metrics like safety. OmniTester [16] extended text-only evaluation to multimodal analysis by generating captions from visualized scenarios using a vision-language model and comparing them with original descriptions from LLM. Notably, a recent survey [17] summarizes existing LLM-based scenario analysis methods. However, none of the papers explicitly assess safety criticality in driving scenarios by using LLMs.
@@ III-C2 Prompt Formulations
The safety-critical metrics prompt follows the structure of $\textit{Prompt}_{\mathrm{cart}}$ and $\textit{Prompt}_{\mathrm{ego}}$, with the LLM acting as a domain expert in collision evaluation. It incorporates CoT reasoning and ICL examples to guide the assessment of overall risk scores for each obstacle, based on thresholds for each safety-critical metric. Our prompt also includes motion descriptions, as shown in Fig. 5(a), focusing on obstacles approaching from the front or sides, since these directions are relevant to evaluate the planner’s performance. This structure enables the LLM to reason about ego-attackers using both quantitative metrics and spatial context.
For trajectory generation, we adopt a different prompting strategy as the LLM transitions from evaluator to generator. The system message instructs the model to synthesize an adversarial trajectory that increases the collision likelihood with the EV, given the scenario context. To enhance the generation performance, we apply Self-Consistency (SC) by prompting the model to produce multiple adversarial trajectory candidates for the same ego-attacker. Each candidate is scored using the safety-critical metrics prompt, and the one with a low-risk score for the ego-attacker is selected.
@@ III-D Structured Output Format
In our LLM-based scenario evaluation and generation framework, we enforce structured output formats using predefined schemes to ensure consistency and reduce hallucination [33]. These formats define fixed keys aligned with specific use cases, including collision evaluation, safety-critical scenario inference, agent-level risk assessment (e.g., risk scores), and trajectory generation parameters (e.g., position). The structure remains flexible and can be adapted to various prompt designs and task requirements.
@@ IV-B Collision Evaluation in Driving Scenarios
This use case evaluates if collisions occur with the EV in driving scenarios. The inputs of our LLM include a map with geometric lanelet information, dynamic states of the EV, and of all traffic participants throughout the simulation’s duration. Our model outputs a boolean flag to indicate if a collision occurs for the EV in the given scenario. We randomly select 100 non-collision and 100 collision scenarios for evaluation from the 6000 total scenarios.
@@ IV-B1 Ablation Study
The performance of our LLM-based evaluator depends on prompt design and advanced techniques with CP, ICL, and CoT reasoning, as described in Section III-B2. To evaluate their impact, an ablation study is conducted for both $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$: Base (no techniques), CP (CP only), CoT (CP + CoT), and All (CP + CoT + ICL).
@@ IV-B2 Comparison Across Different LLMs
We compare the reasoning performance of Gemini-1.5Pro, GPT-4o, and DeepSeek-V3 using the same $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$ and $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ with All configuration applied. As shown in Table II, all models perform significantly better with $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$, confirming the findings in Section IV-B1. Gemini-1.5Pro achieves the highest F1 score (0.21) with $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$, likely due to its larger context window (up to 200k tokens), which helps interpret verbose Cartesian inputs. With ego-centric prompts, DeepSeek-V3 slightly outperforms the others (F1 score 0.90), followed by GPT-4o (0.88) and Gemini-1.5Pro (0.87).
@@ IV-C2 Comparison Across Different LLMs
0022224444666688881010101012121212141414140011112222333344445555Refer to captionTimestepRisk Score (0-5)Original safe scenarioGenerated safety-critical scenario
Figure 8: Risk score of the ego-attacker vehicle’s trajectory, in the original safe scenario and in the generated safety-critical one. A risk score of 5 indicates very low risk, while 0 denotes a collision.
@@ V Conclusion & Outlook
The framework was evaluated in three use cases supported by structured outputs: collision scenario evaluation, safety inference, and safety-critical scenario generation. Across all LLMs, $\mathrm{\textit{Prompt}_{\mathrm{ego}}}$ consistently outperformed $\mathrm{\textit{Prompt}_{\mathrm{cart}}}$, for the Gemini-1.5Pro, achieving F1 scores of 0.87 vs. 0.21 in collision evaluation, and 0.83 vs. 0.62 in safety inference, highlighting the advantage of ego-centric representations for reasoning about risk. In generation, the framework successfully minimized the ego-attacker’s risk score to synthesize safety-critical outcomes. Future work includes extending the framework to 3D simulators such as CARLA, enhancing the generation module to produce more diverse safety-critical scenarios, and using the generated scenarios to further evaluate motion planning algorithms [36].